You are here: Foswiki>Commissioning Web>Logs>21Jan2013_Log (22 Jan 2013, JuanGuerra)Edit Attach

UT 20130121 AO#1

Telescope Operator: Geno.B.(LBTO)
SW Support: Cox,D (Tucson)
AO Support: Jguerra (LBTO)
Telescope Support: Jeff U(LBTO)
Instrument Support: Elliot S (LBTO)


RetroRefrlector on AO#1 (AGW#2). Alignment of the W# with the RR to get the spot back in the IRTC and in the wave front sensor. Pupil wobble analysis. Issues found -fastlink. -Telecope rotator mounted 180 degrees respect the previous mount. -IRTC controller OFF. -IRS server stop. -adscec dx disk not mount


The W unit was aligned to the RR and a new HEXAPOD POSITION was found: x:-0.91 y:1.46 z:3.65 Rx:-359 Ry:-48, Rz:0

The sistem could be close the loop at bin 1 without any problem.

The check of the pupil wobble taking data of camara lens position, rotator position and thin shell offload. In the Arcetri daily report will give a more detailed information

-- JuanGuerra - 21 Jan 2013
Topic revision: r2 - 22 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback